OpenStudy (anonymous):

When -4(2x^2-3x-12) is simplified, what is the value of the constant term? A. -8 B.12 C.48 D.52

OpenStudy (mysangels):

Follow the Order of Operations Parentheses Exponents Multiplication Division Addition Subtraction

OpenStudy (jdoe0001):

distribute the -4, see what you get

OpenStudy (anonymous):

I got -32x-12x-48 Do I add -32 and -12x

OpenStudy (mysangels):

If they are like terms, yes.

OpenStudy (jdoe0001):

-32x?

OpenStudy (anonymous):

Yes I go that

OpenStudy (jdoe0001):

\(\bf {\color{brown}{ -4}}(2x^2-3x-12)\implies ({\color{brown}{ -4}}\cdot 2x^2)+({\color{brown}{ -4}}\cdot -3x)+({\color{brown}{ -4}}-12)\)

OpenStudy (jdoe0001):

hmm missed a multiplication bit \(\bf {\color{brown}{ -4}}(2x^2-3x-12)\implies ({\color{brown}{ -4}}\cdot 2x^2)+({\color{brown}{ -4}}\cdot -3x)+({\color{brown}{ -4}}\cdot -12)\) but anyhow, you wouldn't get a 32x either way, and the in the "constant" is the one without a variable

OpenStudy (anonymous):

The answer is d?

OpenStudy (jdoe0001):

hmm dunno... what did you get for the constant?

OpenStudy (anonymous):

Well I did what you told me to do in the last status and I got. 52

OpenStudy (anonymous):

What would be the constant term? Well what is a constant term?

OpenStudy (mysangels):

a constant term is a number with no variable, which is why it is constant with no variable that can be any number in the universe

OpenStudy (anonymous):

I tnk the answer is d cause that's what I got as my final answer

OpenStudy (anonymous):

??